Audrey has 28 dimes and quarters in her piggy bank with a total value of $4.75. How many of the coins are dimes and how many are quarters?: Audrey has 28 dimes and quarters in her piggy bank with a total value of $4.75. How many of the coins are dimes and how many are quarters?
Answer by Fombitz(1756)

Let's call dimes, D, and quarters, Q.
1.D+Q=28 Total number of coins.
2.0.10*D+0.25*Q=4.75 Total value of coins.
Let's multiply equation 2 by 100 to get rid of decimals.
1.D+Q=28
2.10D+25Q=475
Use eq. 1 to solve for D and a function of Q.
1.D+Q=28
D=28-Q
Now substitute this expression for D in eq. 2,
2.10D+25Q=475
10(28-Q)+25Q=475
Solve for Q.
280-10Q+25Q=475
280+15Q=475
15Q=195
Q=13
Use eq.1 to solve for D.
D=28-Q
D=28-13
D=15
Audrey has 15 dimes and 13 quarters.